IN this example the soil provides an average thrust T of 2.0 x 106 N and acts about a third of the way up the wall. The weight acts 1.0 m from the right hand corner of the wall. Calculate the weight of wall just needed for the wall to be stable and not topple over.
1
Expert's answer
2020-09-28T06:03:40-0400

"\Sigma M_A=0\to 3\cdot T=1\cdot W"


"W=3T=3\cdot2\cdot10^6=6\cdot10^6(N)"